# reaction term and pure phase initial datum

### v1

@@ 4 Uniqueness of weak solutions
First, we set $\varphi=\varphi_{2}-\varphi_{1}$ and consider the weak formulation for the difference that reads as
$\mathopen{\langle}\partial_{t}\varphi,v\mathclose{\rangle}_{V}+\int_{\Omega}\nabla\mu\cdot\nabla v=-\int_{\Omega}m(\cdot)\varphi v+\int_{\Omega}(h(\cdot,\varphi_{2},g)-h(\cdot,\varphi_{1},g))v,$ (4.1)
@@
$\frac{d}{dt}\varphi_{\Omega}=-(m(\cdot)\varphi)_{\Omega}+(h(\cdot,\varphi_{2},g)-h(\cdot,\varphi_{1},g))_{\Omega}.$ (4.2)
Multiplying now the above equation by $v=\varphi_{\Omega}+\mathop{\rm sign}\nolimits\varphi_{\Omega}$ and using the Lipschitz continuity of $h$, we obtain
$\frac{d}{dt}\Big{(}\frac{1}{2}|{\varphi_{\Omega}}|^{2}+|\varphi_{\Omega}|\Big{)}\leq C(|\varphi_{\Omega}|^{2}+|\varphi_{\Omega}|+\left\|\varphi\right\|^{2}).$ (4.3)
@@
$\quad=-\int_{\Omega}m(\cdot)\varphi{\cal N}(\varphi-\varphi_{\Omega})$
$\qquad+\int_{\Omega}\Big{(}(h(\cdot,\varphi_{2},g)-h(\cdot,\varphi_{1},g))-((h(\cdot,\varphi_{2},g)-h(\cdot,\varphi_{1},g)))_{\Omega}\Big{)}{\cal N}(\varphi-\varphi_{\Omega})$
$\qquad-\int_{\Omega}(\pi(\varphi_{2})-\pi(\varphi_{1}))(\varphi-\varphi_{\Omega})-\int_{\Omega}(\beta(\varphi_{2})-\beta(\varphi_{1}))\varphi_{\Omega}=:\sum_{i=1}^{4}I_{i}.$ (4.4)
@@
$I_{1}+I_{2}+I_{3}$ $\leq C\left\|\varphi\right\|^{2}+C\left\|\varphi-\varphi_{\Omega}\right\|^{2}_{*}\leq\delta\left\|\nabla\varphi\right\|^{2}+C_{\delta}\left\|\varphi-\varphi_{\Omega}\right\|^{2}_{*}+C|\varphi_{\Omega}|^{2},$
$I_{4}$ $\leq C(\left\|\beta(\varphi_{2})\right\|_{1}+\left\|\beta(\varphi_{1})\right\|_{1})|\varphi_{\Omega}|.$
@@
Here, we just emphasize the key differences with respect to the previous proof. As above, we set $\varphi=\varphi_{2}-\varphi_{1}$ and observe that
$\mathopen{\langle}\partial_{t}\varphi,v\mathclose{\rangle}_{V}+\int_{\Omega}\nabla\mu\cdot\nabla v=-\int_{\Omega}m(\cdot)\varphi v+\int_{\Omega}(h(\cdot,\varphi_{2},g)-h(\cdot,\varphi_{1},g))v,$
@@
$\quad-\int_{\Omega}m(\cdot)\varphi{\cal N}(\varphi-\varphi_{\Omega})$
$\quad+\int_{\Omega}\Big{(}(h(\cdot,\varphi_{2},g)-h(\cdot,\varphi_{1},g))-((h(\cdot,\varphi_{2},g)-h(\cdot,\varphi_{1},g)))_{\Omega}\Big{)}{\cal N}(\varphi-\varphi_{\Omega})$
$\quad-\int_{\Omega}(\pi(\varphi_{2})-\pi(\varphi_{1}))(\varphi-\varphi_{\Omega})-\int_{\Omega}(\beta(\varphi_{2})-\beta(\varphi_{1}))\varphi_{\Omega},$
where the second term on the left-hand side arises from $\int_{\Omega}(a\varphi+\beta(\varphi_{2})-\beta(\varphi_{1}))\varphi$ using A1, A4, and (2.22). Observe now that, using the Young inequality, we have
@@ 5 Application to tumor growth models
With no claim to be exhaustive, let us briefly review some related works dealing with singular potentials like (1.2). For those models, a crucial point is to handle the mass evolution; in particular one would avoid inconsistency of the model and have non physical solutions $\varphi$ that satisfy $|(\varphi(t))_{\Omega}|>1$ for some time $t\in[0,T]$. To overcame this possible situation one usually proceed as we did and requires that the source term $S$, which rules the mass evolution of the order parameter, is given by (5.6). In this direction, we refer to [39, 54, 24]. Alternatively, some authors suggest to introduce some regularization terms in the system to allow for more general potentials and source terms (see, for instance, [12, 55] and the references therein). Finally, let us point out that nonlocal models in the context of phase field tumor growth models are also investigated as well (see, e.g., [55, 25]).
@@
and assume that $\mathbbm{h}$ is uniformly bounded, $\chi$ is a nonnegative constant, and $n_{B}\in L^{\infty}(Q)$ with $0\leq n_{B}\leq 1$ a.e. in $Q$. Then, problem (5.2)–(5.4) with $j=1$ admits a weak solution $(\varphi,\mu,n)$ such that
@@
and assume that $\mathbbm{h}$ is uniformly bounded, $\chi$ is a nonnegative constant, and $n_{B}\in L^{\infty}(Q)$ with $0\leq n_{B}\leq 1$ a.e. in $Q$. Then, problem (5.2)–(5.5) with $j=2$ admits a weak solution $(\varphi,\mu,n)$ such that
@@
If $\mathbbm{h}$ satisfies an assumption like (2.22), then continuous dependence estimates can be proven arguing along the lines of Theorem 2.8 and Theorem 2.9. In this case, both the local and the nonlocal problem are well posed.
@@ 6 Final comments and future work
The present results are a first step towards a more general characterization of source terms which allow a pure initial datum. Another interesting, and perhaps nontrivial, issue is the extension of our results to other models. For instance, Cahn–Hilliard–Navier–Stokes system with sources (see [7, 42] and references therein), fractional Cahn–Hilliard equations with sources (see, for instance, [27, 29] for the conserved case), or multi-component Cahn–Hilliard equations with sources (see [17], see also [28] for the general conserved case).
@@ Appendix A Appendix
Property (P1) follows from [35, Sec. 5] and [52]. Let us focus only on the proof of property (P2). Since $\widehat{\beta}$ is strictly convex with $\widehat{\beta}(0)=0$, one has that $\beta$ is increasing, $\beta(0)=0$, with $\beta<0$ in $(-1,0)$, and $\beta>0$ in $(0,1)$.
@@
Acknowledgements. The authors are members of Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA), Istituto Nazionale di Alta Matematica (INdAM). Besides, this research has been performed in the framework of the MIUR-PRIN Grant 2020F3NCPX “Mathematics for industry 4.0 (Math4I4)”. The present research has been supported by MUR, grant Dipartimento di Eccellenza 2023-2027.

### v2

@@ 4 Uniqueness of weak solutions
First, we set $\varphi=\varphi_{2}-\varphi_{1}$ and consider the weak formulation for the difference that reads, recalling that $h$ is constant, as
$\mathopen{\langle}\partial_{t}\varphi,v\mathclose{\rangle}_{V}+\int_{\Omega}\nabla\mu\cdot\nabla v=-\int_{\Omega}m(\cdot)\varphi v{,}$ (4.1)
@@
$\frac{d}{dt}\varphi_{\Omega}=-(m(\cdot)\varphi)_{\Omega}{.}$ (4.2)
Multiplying now the above equation by $v=\varphi_{\Omega}+\mathop{\rm sign}\nolimits\varphi_{\Omega}$, we obtain
$\frac{d}{dt}\Big{(}\frac{1}{2}|{\varphi_{\Omega}}|^{2}+|\varphi_{\Omega}|\Big{)}\leq C(|\varphi_{\Omega}|^{2}{+|\varphi_{\Omega}|}).$ (4.3)
@@
$\quad=-\int_{\Omega}m(\cdot)\varphi{\cal N}(\varphi-\varphi_{\Omega})-\int_{\Omega}(\pi(\varphi_{2})-\pi(\varphi_{1}))(\varphi-\varphi_{\Omega})$
$\qquad-\int_{\Omega}(\beta(\varphi_{2})-\beta(\varphi_{1}))\varphi_{\Omega}=:\sum_{i=1}^{3}I_{i}.$ (4.4)
@@
$I_{1}+{I_{2}}$ $\leq C\left\|\varphi\right\|^{2}+C\left\|\varphi-\varphi_{\Omega}\right\|^{2}_{*}\leq\delta\left\|\nabla\varphi\right\|^{2}+C_{\delta}\left\|\varphi-\varphi_{\Omega}\right\|^{2}_{*}+C|\varphi_{\Omega}|^{2},$
${I_{3}}$ $\leq C(\left\|\beta(\varphi_{2})\right\|_{1}+\left\|\beta(\varphi_{1})\right\|_{1})|\varphi_{\Omega}|.$
@@
Here, we just emphasize the key differences with respect to the previous proof. Also in this case, setting $\varphi=\varphi_{2}-\varphi_{1}$, we have (cf. (4.1))
$\mathopen{\langle}\partial_{t}\varphi,v\mathclose{\rangle}_{V}+\int_{\Omega}\nabla\mu\cdot\nabla v=-\int_{\Omega}m(\cdot)\varphi v{,}$
@@
$\quad-\int_{\Omega}m(\cdot)\varphi{\cal N}(\varphi-\varphi_{\Omega}){-\int_{\Omega}(\pi(\varphi_{2})-\pi(\varphi_{1}))(\varphi-\varphi_{\Omega})}$
$\quad-\int_{\Omega}(\beta(\varphi_{2})-\beta(\varphi_{1}))\varphi_{\Omega},$
where the second term on the left-hand side arises from $\int_{\Omega}(a\varphi+\beta(\varphi_{2})-\beta(\varphi_{1}))\varphi$ using A1, and A4. Observe now that, using the Young inequality, we have
@@ 5 Application to tumor growth models
With no claim to be exhaustive, let us briefly review some related works dealing with singular potentials like (1.2). For those models, a crucial point is to handle the mass evolution; in particular one would avoid inconsistency of the model and have non physical solutions $\varphi$ that satisfy $|(\varphi(t))_{\Omega}|>1$ for some time $t\in[0,T]$. To overcome this possible situation one usually proceed as we did and requires that the source term $S$, which rules the mass evolution of the order parameter, is given by (5.6). In this direction, we refer to [40, 56, 25]. Alternatively, some authors suggest to introduce some regularization terms in the system to allow for more general potentials and source terms (see, for instance, [12, 57] and the references therein). Finally, let us point out that nonlocal models in the context of phase field tumor growth models are also investigated as well (see, e.g., [57, 26]).
@@
and assume that $\mathbbm{h}$ is continuous and uniformly bounded, $\chi$ is a nonnegative constant, and $n_{B}\in L^{\infty}(Q)$ with $0\leq n_{B}\leq 1$ a.e. in $Q$. Then, problem (5.2)–(5.4) with $j=1$ admits a weak solution $(\varphi,\mu,n)$ such that
@@
and assume that $\mathbbm{h}$ is continuous and uniformly bounded, $\chi$ is a nonnegative constant, and $n_{B}\in L^{\infty}(Q)$ with $0\leq n_{B}\leq 1$ a.e. in $Q$. Then, problem (5.2)–(5.5) with $j=2$ admits a weak solution $(\varphi,\mu,n)$ such that
@@
If the function $\mathbbm{h}$ does not depend on the phase variable $\varphi$, then continuous dependence estimates can be proven arguing along the lines of Theorem 2.8 and Theorem 2.9. In this case, both the local and the nonlocal problem are well posed.
@@ 6 Final comments and future work
The present results are a first step towards a more general characterization of source terms which allow a pure initial datum. Another interesting, and perhaps nontrivial, issue is the extension of our results to other models. For instance, Cahn–Hilliard–Navier–Stokes system with sources (see [43] and references therein), fractional Cahn–Hilliard equations with sources (see, for instance, [28, 30] for the conserved case), or multi-component Cahn–Hilliard equations with sources (see [17], see also [29] for the general conserved case).
@@ Appendix A Appendix
Property (P1) follows from [36, Sec. 5] and [54]. Let us focus only on the proof of property (P2). Since $\widehat{\beta}$ is strictly convex with $\widehat{\beta}(0)=0$, one has that $\beta$ is increasing, $\beta(0)=0$, with $\beta<0$ in $(-1,0)$, and $\beta>0$ in $(0,1)$.
@@
Acknowledgements. The authors thank the reviewers for their valuable comments and suggestions. The authors are members of Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA), Istituto Nazionale di Alta Matematica (INdAM). Besides, this research has been performed in the framework of the MIUR-PRIN Grant 2020F3NCPX “Mathematics for industry 4.0 (Math4I4)”. AS gratefully acknowledge some support from the Alexander von Humboldt Foundation. The present research has been supported by MUR, grant Dipartimento di Eccellenza 2023-2027.